# Tuning the spin dynamics and magnetic phase transitions of the Cantor alloy via composition and sample processing protocols: A muon spin relaxation study

### v1

@@ III.1 SQUID Magnetometry
Figure 1: Temperature-dependent molar susceptibility for the five Cantor alloys studied in this work. Measurements were performed in a zero-field-cooled warming sequence in an applied field of 100 Oe. Note that the cold-worked and Mn0 data have been scaled down by factors of 0.5 and 0.0004, respectively, for convenience.
@@ III.2 Zero-Field $\mu$SR Data
Figure 2: Representative spectra for each Cantor alloy sample (labeled on the individual plots). Filled symbols show the experimental data, and the solid black curves are fits using a stretched exponential function (a-e) or a Kubo-Toyabe function (f). Note the shorter time window visible in panel (f), in contrast to the longer time window used in the other panels.
@@
The temperature dependence of the relaxation rate $\lambda$ is shown in Fig. 4(b). A prominent peak centered around 82 K appears for the pristine and premag samples and 78 K for Mn0. The temperatures around which the peaks are centered correspond well to the transition temperatures already identified for each sample. This type of peak in the relaxation rate is evidence of “critical slowing down” of the spin dynamics in the vicinity of the transition, as observed in canonical spin glasses and continuous phase transitions [22, 23]. The cold-worked and Mn30 samples also exhibit a peak in $\lambda$ centered around 43 K and 186 K, respectively, but the peak is much broader for the cold-worked sample and much smaller in magnitude for Mn30. The broadness in the cold-worked sample is attributable to the wide spread of transition temperatures throughout the sample volume, while the smaller peak observed for Mn30 reveals a suppression of the critical spin dynamics in this sample compared to the others.
The $\beta$ values extracted from the ZF fits are shown in Fig. 4(c). For all samples, the values at high temperature are significantly above unity, consistent with the somewhat Gaussian-like curvature observed in Fig. 2. With decreasing temperature, $\beta$ decreases until it becomes approximately constant below the transition temperature. In our initial fits, we allowed $\beta$ to vary independently for each run, and we observed that the values naturally converged to approximately 0.5 for the pristine, premag, and cold-worked samples, 0.75 for Mn30, and 1.0 for Mn0, with some scatter in the results on the order of 10%. For simplicity, we then conducted another set of fits with $\beta$ fixed to these values below the transition. The value of $\beta$ is physically significant: in situations with a single relaxation channel with just one relaxation rate, $\beta$ is expected to be unity, while the effect of multiple relaxation channels and/or a distribution of relaxation rates is typically modeled with $\beta<1$ [24, 25, 26]. The sub-exponential behavior observed in the equiatomic and Mn30 samples is similar to the behavior seen in a five-component high-entropy oxide [27], suggesting this property may be a general feature of $\mu$SR spectra in magnetic high-entropy materials. The systematic increase of the low-temperature $\beta$ value from 0.5 in the equiatomic samples to 0.75 in Mn30 to 1.0 in Mn0 further supports the notion that increased configurational entropy drives the sub-exponential relaxation.
@@ III.3 Longitudinal-Field $\mu$SR Data
Figure 5: Representative LF $\mu$SR spectra for the pristine, Mn0, and Mn30 samples. Filled symbols represent the experimental data, the black curves the fits using a stretched exponential function.
@@
The long-time relaxation rate $\lambda_{LF}$ is displayed in Fig. 6(b). The pristine sample features a large, sharp peak centered around 82 K, confirming that a critical slowing down of the spin fluctuations occurs in the vicinity of the transition temperature. Mn30 likewise shows a sharp peak centered around 183 K, although the magnitude is much smaller than that of the pristine sample. In contrast, Mn0 shows a broad, low peak centered around 97 K. The breadth of the peak is due to the smeared out transition in an applied field, while the diminished height of the peak indicates that the critical slowing down observed in the pristine sample and (to a slightly lesser extent) the Mn30 sample is greatly suppressed in the Mn0 sample. These aspects of the peak shape differ from the ZF data for Mn0, which exhibits a sharp peak in the relaxation rate at the transition temperature, suggesting the application of a relatively modest field of 0.1 T significantly modifies the internal spin dynamics.
@@
where $\Delta=\gamma_{\mu}B_{i}$ is the product of the muon’s gyromagnetic ratio $\gamma_{\mu}$ and the root mean square internal field $B_{i}$, and $w_{L}=\gamma_{\mu}B_{L}$ is the Larmor frequency of the muon spin in the longitudinal field $B_{L}$. As has been done elsewhere [27], $\Delta$ can be estimated from the Kubo-Toyabe relaxation rate observed at low temperature in Mn0, which is $120\pm 10$ $\mu$s-1. We note that Eq. 3 relies on the assumption that the asymmetry exhibits purely exponential relaxation in the paramagnetic phase, which is only approximately the case in the present study, but this calculation nevertheless yields a reasonable estimate of the spin fluctuation rate as a function of temperature. The calculated spin fluctuation rate is shown for the pristine, Mn0 and Mn30 samples in Fig. 7.
@@ Acknowledgements
We thank the staff at TRIUMF for their valuable help and support during the $\mu$SR experiments, along with Raju Baral, Braedon Jones, Kane Fanning, Karolyn Mocellin, Christiana Suggs, and Parker Hamilton for assistance with the $\mu$SR data collection. E.Z. and B.A.F. acknowledge the College of Physical and Mathematical Sciences at Brigham Young University for funding support. T.A.E. and M.W.M. acknowledge support from NSF No. DMR-1708410. Synthesis and characterization facilities at the University of Florida were developed under support from NSF-CAREER 1453752 (J.J.H.). RD is supported by the Center for Integrated Nanotechnologies, an Office of Science user facility operated for the U.S. Department of Energy. This article has been authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all right, title, and interest in and to the article and is solely responsible for its contents. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan .

### v2

@@ III Results and Analysis
The new results and analysis presented in the following are interspersed with a selection of results published previously [15, 16] to facilitate comparison of different samples and highlight emerging trends. New in this work are magnetometry data on the premag sample; ZF $\mu$SR data on the cold-worked and premag samples; LF $\mu$SR data on the pristine sample; analysis of the early-time and time-integrated ZF asymmetry for all samples; and extraction of the spin fluctuation rate and spin-flip activation energy from detailed analysis of the LF data on the pristine, Mn0, and Mn30 samples. Results carried over from previous publications include magnetometry data for all but the premag sample; ZF $\mu$SR data and initial anlaysis for the pristine, Mn0, and Mn30 samples; and LF $\mu$SR relaxation rates for Mn0 and Mn30. The present work builds upon the previously published $\mu$SR results [15, 16] by focusing on the microscopic details of the magnetic transition and dynamics, including homogeneity, spin fluctuation rates and energies, and the tunability of these properties via composition, sample processing, and magnetic field.
@@ III.1 SQUID Magnetometry
Figure 1: Temperature-dependent molar susceptibility for the five Cantor alloys studied in this work. Measurements were performed in a zero-field-cooled warming sequence in an applied field of 100 Oe. Note that the cold-worked and Mn0 data have been scaled down by factors of 0.5 and 0.0004, respectively, for convenience. Data for the pristine, Mn30, Mn0, and cold-worked samples published previously in [15, 16].
@@ III.2 Zero-Field $\mu$SR Data
Figure 2: Representative spectra for each Cantor alloy sample (labeled on the individual plots). Filled symbols show the experimental data, and the solid black curves are fits using a stretched exponential function (a-e) or a Kubo-Toyabe function (f). Note the shorter time window visible in panel (f), in contrast to the longer time window used in the other panels. Panels (a), (c), and (e) contain data published previously [15, 16].
@@
The temperature dependence of the relaxation rate $\lambda$ is shown in Fig. 4(b). A prominent peak centered around 82 K appears for the pristine and premag samples and 78 K for Mn0. The temperatures around which the peaks are centered correspond well to the transition temperatures already identified for each sample. In the fast fluctuation limit, $\lambda$ is inversely proportional to the spin fluctuation rate [21], so an increase in $\lambda$ denotes a decrease in fluctuation rate. Thus, the peak observed in the relaxation rate is evidence of “critical slowing down” of the spin dynamics in the vicinity of the transition, as seen in canonical spin glasses and continuous phase transitions [22, 21]. The cold-worked and Mn30 samples also exhibit a peak in $\lambda$ centered around 43 K and 186 K, respectively, but the peak is much broader for the cold-worked sample and much smaller in magnitude for Mn30. The broadness in the cold-worked sample is attributable to the wide spread of transition temperatures throughout the sample volume, while the smaller peak observed for Mn30 reveals a suppression of the critical spin dynamics in this sample compared to the others.
The $\beta$ values extracted from the ZF fits are shown in Fig. 4(c). For all samples, the values at high temperature are significantly above unity, consistent with the somewhat Gaussian-like curvature observed in Fig. 2. With decreasing temperature, $\beta$ decreases until it becomes approximately constant below the transition temperature. In our initial fits, we allowed $\beta$ to vary independently for each run, and we observed that the values naturally converged to approximately 0.5 for the pristine, premag, and cold-worked samples, 0.75 for Mn30, and 1.0 for Mn0, with some scatter in the results on the order of 10%. For simplicity, we then conducted another set of fits with $\beta$ fixed to these values below the transition. The value of $\beta$ is physically significant: in situations with a single relaxation channel with just one relaxation rate, $\beta$ is expected to be unity, while the effect of multiple relaxation channels and/or a distribution of relaxation rates is typically modeled with $\beta<1$ [21, 23, 24]. The sub-exponential behavior observed in the equiatomic and Mn30 samples is similar to the behavior seen in a five-component high-entropy oxide [25], suggesting this property may be a general feature of $\mu$SR spectra in magnetic high-entropy materials. The configurational entropy is maximized for the equiatomic composition and decreases as the composition moves further away from equal proportions. As such, the systematic increase of the low-temperature $\beta$ value from 0.5 in the equiatomic samples to 0.75 in Mn30 to 1.0 in Mn0 further supports the notion that increased configurational entropy drives the sub-exponential relaxation.
@@ III.3 Longitudinal-Field $\mu$SR Data
Figure 5: Representative LF $\mu$SR spectra for the pristine, Mn0, and Mn30 samples for an applied 0.1 T field. Filled symbols represent the experimental data, the black curves the fits using a stretched exponential function.
@@
The long-time relaxation rate $\lambda_{LF}$ is displayed in Fig. 6(b). The pristine sample features a large, sharp peak centered around 82 K, confirming that a critical slowing down of the spin fluctuations occurs in the vicinity of the transition temperature. Mn30 likewise shows a sharp peak centered around 183 K, although the magnitude is much smaller than that of the pristine sample. In contrast, Mn0 shows a broad, low peak centered around 97 K. The breadth of the peak is due to the smeared out transition in an applied field, while the diminished height of the LF peak indicates that the critical slowing down observed in the pristine sample and (to a slightly lesser extent) the Mn30 sample is greatly suppressed in the Mn0 sample. These aspects of the peak shape differ from the ZF data for Mn0, which exhibits a sharp peak in the relaxation rate at the transition temperature, suggesting the application of a relatively modest field of 0.1 T significantly modifies the internal spin dynamics. We note that the shift of the peak center from 78 K in ZF to 97 K in LF is due to ferromagnetic nature of Mn0, in contrast to the other compositions, which do not show a large change in the temperature of the peak center with applied field.
@@
where $\Delta=\gamma_{\mu}B_{i}$ is the product of the muon’s gyromagnetic ratio $\gamma_{\mu}$ and the root mean square internal field $B_{i}$, and $\omega_{L}=\gamma_{\mu}B_{L}$ is the Larmor frequency of the muon spin in the longitudinal field $B_{L}$. As has been done elsewhere [25], $\Delta$ can be estimated from the Kubo-Toyabe relaxation rate observed at low temperature, which is $120\pm 10$ $\mu$s-1 for Mn0. Since the pristine and Mn30 samples do not exhibit a clear Kubo-Toyabe relaxation pattern, we use the value of $\Delta$ determined from the Mn0 data as an upper bound when calculating $\nu$. We base this on the fact that the net effective moment increases as the Mn content is decreased [16], leading to a larger expected local field for Mn0. Thus, the calculated values of $\nu$ for the pristine and Mn30 samples should be considered upper bounds. We further note that Eq. 3 relies on the assumption that the asymmetry exhibits purely exponential relaxation in the paramagnetic phase, which is only approximately the case in the present study, but this calculation nevertheless yields a reasonable estimate of the spin fluctuation rate as a function of temperature. The calculated spin fluctuation rate is shown for the pristine, Mn0 and Mn30 samples in Fig. 7.
@@ Acknowledgements
We thank the staff at TRIUMF for their valuable help and support during the $\mu$SR experiments, along with Raju Baral, Braedon Jones, Kane Fanning, Karolyn Mocellin, Christiana Suggs, and Parker Hamilton for assistance with the $\mu$SR data collection. E.Z. and B.A.F. acknowledge the College of Physical and Mathematical Sciences at Brigham Young University for funding support. T.A.E. and M.W.M. acknowledge support from NSF No. DMR-1708410. Synthesis and characterization facilities at the University of Florida were developed under support from NSF-CAREER 1453752 (J.J.H.). RD is supported by the Center for Integrated Nanotechnologies, an Office of Science user facility operated for the U.S. Department of Energy. This article has been authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan .